OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!! Please help me to simplify this rational expression - x^2-1 x-1 -------- / ----- 2x 4x^2 **Note the / means to divide

OpenStudy (mertsj):

Why did you repost this? I showed you how to do it. Even factored it for you.

OpenStudy (abbycross167):

Sir/ma'am? I reposted this question because no one was answering me and the only person I that was helping me went offline...

OpenStudy (mertsj):

\[\frac{(x-1)(x+1)}{2x}\div \frac{x-1}{4x^2} = \]

OpenStudy (mertsj):

\[\frac{(x-1)(x+1)}{2x} \times \frac{4x^2}{x-1}\]

OpenStudy (mertsj):

Do you understand so far?

OpenStudy (abbycross167):

Yes sir/ma'am but how did you get the (x-1)(x-1)?

OpenStudy (abbycross167):

@Mertsj

OpenStudy (mertsj):

It's (x-1)(x+1) . Those are the factors of x^2-1